Three athletes are tugging on a ring. Martha pulls with a force of 150N north. Evelyn pulls with a force of 150N directly east. Sara pulls with a force of 350N directly southeast (135degrees clockwise from north). What is the resultant force from the three forces. Please show step by step and all work.Three athletes are tugging on a ring. Martha pulls with a force of 150N north. Evelyn pulls with a force of 150N directly east. Sara pulls with a force of 350N directly southeast (135degrees clockwise from north). What is the resultant force from the three forces. Please show step by step and all work.
Answer #1

Let East be +x axis


Let

F1 = 150 N

F2 = 150 N

F3 = 350 N

Fnetx = F1*cos(90) + F2*cos(0) + F3*cos(45)

= 0 + 150 + 350*cos(45)

= 397.5 N

Fnety = F1*cos(0) + F2*cos(90) - F3*sin(45)

= 0 + 150 - 350*sin(45)

= -97.5 N

Fnet = sqrt(Fnetx^2 + Fnety^2)

= sqrt(397.5^2 + 97.5^2)

= 409.3 N <<<<<<<<<<<<<<<<<<<<<<---------Answer